### v1

# Sketch-Guided Constrained Decoding for Boosting Blackbox Large Language Models without Logit Access

Abstract
Constrained decoding, a technique for enforcing constraints on language model outputs, offers a way to control text generation without retraining or architectural modifications. Its application is, however, typically restricted to models that give users access to next-token distributions (usually via softmax logits), which poses a limitation with blackbox large language models (LLMs). This paper introduces sketch-guided constrained decoding (SGCD), a novel approach to constrained decoding for blackbox LLMs, which operates without access to the logits of the blackbox LLM. SGCD utilizes a locally hosted auxiliary model to refine the output of an unconstrained blackbox LLM, effectively treating this initial output as a “sketch” for further elaboration. This approach is complementary to traditional logit-based techniques and enables the application of constrained decoding in settings where full model transparency is unavailable. We demonstrate the efficacy of SGCD through experiments in closed information extraction and constituency parsing, showing how it enhances the utility and flexibility of blackbox LLMs for complex NLP tasks. [^1] We will release the code and data upon publication.
1 Introduction
Figure 1: Overview of sketch-guided constrained decoding (SGCD). In the initial sketching phase, a blackbox LLM generates a preliminary “sketch” answer without applying any constraints. Then, in the constrained generation phase, an auxiliary model, the constrained generator, refines the sketch. The refined, final output respects the specified constraints by construction.
Large language models (LLMs) have seen a remarkable expansion in scope, being used for diverse tasks including tool interaction, SQL translation, robotic navigation and item recommendations, where adherence to specific constraints is paramount (Bubeck et al., 2023; Schick et al., 2023; Poesia et al., 2022; Shah et al., 2022; Zhang et al., 2023; Hua et al., 2023). Despite their versatility, LLMs often struggle with constraint adherence in few-shot scenarios, leading to outputs that violate task-specific requirements (Chen and Wan, 2023; Agrawal et al., 2023; Huang et al., 2023).
Constrained decoding offers a solution, restricting model outputs to respect predefined constraints without necessitating model retraining or architectural modifications (Poesia et al., 2022; Shin et al., 2021; Beurer-Kellner et al., 2023; Scholak et al., 2021; Geng et al., 2023). However, existing constrained decoding methods require access to the model’s logits during inference, which is not always feasible in practice (cf. Appendix A). Since the most powerful LLMs tend to be commercial and blackbox (Lee et al., 2023), this has restricted the application of constrained decoding methods.
Contributions. In response, we present sketch-guided constrained decoding (SGCD), which bypasses the need for direct logit access. SGCD uses a locally hosted auxiliary model to refine the outputs of a blackbox LLM while respecting the specified constraints. We validate our method on closed information extraction and constituency parsing, two challenging NLP tasks that require the model to generate highly structured and grounded outputs. Our experiments show that SGCD significantly boosts the performance of LLMs and beats previous approaches by a wide margin.
2 Method
SGCD splits the constrained decoding task into two distinct phases: sketching and constrained generation.
During sketching, a sketcher—a powerful blackbox LLM denoted as $P_{\text{sk}}$—is employed. It interprets an instruction $I$ alongside a set of demonstration pairs $D=\{(x^{i},y^{i})\}_{i=1}^{n}$, producing a preliminary draft $y^{*}$ via unconstrained decoding:
$y^{*}\approx\operatorname*{arg\,max}_{y\in S}P_{\text{sk}}(y\,|\,I,D,x),$ (1)
where $S$ is the set of all possible sequences.
Constrained generation is done by a constrained generator, a smaller-scale, locally hosted LLM $P_{\text{cg}}$. Given an instruction $I_{\text{cg}}$, a set of input–sketch–output demonstrations $D_{\text{cg}}=\{(x^{i},y^{i},z^{i})\}_{i=1}^{n}$, the original input $x$, and the sketch $y^{*}$, it refines $y^{*}$ into
$z^{*}\approx\operatorname*{arg\,max}_{z\in S\cap C}P_{\text{cg}}(z\,|\,I_{\text{cg}},D_{\text{cg}},x,y^{*}),$ (2)
subject to constraints $C$. (Optionally, $x$ and $x^{i}$ may be omitted, with loss of information.)
The sketcher’s output $y^{*}$ is typically of high quality, encapsulating the necessary information for the constrained generator to produce the final sequence $z^{*}$ that adheres to the constraints $C$. Given the quality of $y^{*}$, the constrained generator can be implemented using a much smaller model, as its primary task is to rewrite the sketch $y^{*}$ with the help of constrained decoding, thus facilitating deployment on standard consumer-grade hardware.
On the contrary, classical, direct few-shot prompting with constrained decoding would usually require a larger constrained generator $P_{\text{cg}}$ to be run locally, in order to find
$w^{*}\approx\operatorname*{arg\,max}_{w\in S\cap C}P_{\text{cg}}(w\,|\,I,D,x).$ (3)
Another basic alternative, unconstrained few-shot prompting (Brown et al., 2020), yields $y^{*}$ as the end product.
SGCD builds on the expectation that the constrained refined output $z^{*}$ should be at least as good as both $y^{*}$ (as $z^{*}$ respects the constraints) and $w^{*}$ (as $P_{\text{sk}}$ is a more powerful LLM than $P_{\text{cg}}$).
3 Experiments
Wiki-NRE SynthIE-text
Precision Recall F1 Precision Recall F1
With logit access
     LLaMA-2-7B + CD 31.7± 2.7 36.8± 2.9 31.9± 2.5 36.6± 2.3 29.4± 2.0 31.2± 2.0
     LLaMA-2-13B + CD 31.4± 2.6 40.7± 3.0 33.1± 2.5 38.9± 2.4 32.6± 2.1 33.6± 2.1
Without logit access
     GPT4 43.2± 2.7 47.2±3.0 42.8± 2.6 46.8± 2.3 45.3± 2.2 45.8± 2.2
     GPT4 + SGCD 56.9± 3.2 (normal-$\uparrow$13.7) 47.4± 2.9 (normal-$\uparrow$2.1) 48.6± 2.8 (normal-$\uparrow$2.8) 53.2± 7.9 (normal-$\uparrow$3.5) 43.0± 2.3 (normal-$\downarrow$-2.3) 45.6± 2.2 (normal-$\downarrow$-0.2)
     GPT3.5 29.6± 2.7 27.0± 2.7 26.7± 2.5 25.7± 2.0 24.0± 1.9 24.6± 1.9
     GPT3.5 + SGCD 47.6± 3.3 (normal-$\uparrow$18.0) 37.9± 2.8 (normal-$\uparrow$10.9) 39.4± 2.8 (normal-$\uparrow$12.7) 44.2± 2.8 (normal-$\uparrow$18.5) 32.4± 2.1 (normal-$\uparrow$8.4) 35.3± 2.1 (normal-$\uparrow$10.7)
     Claude 37.6± 3.1 31.3± 2.8 32.5± 2.7 27.5± 2.0 27.6± 2.0 27.4± 2.0
     Claude + SGCD 49.4± 2.5 (normal-$\uparrow$11.8) 37.5± 2.8 (normal-$\uparrow$6.2) 40.3± 2.9 (normal-$\uparrow$7.8) 46.2± 2.5 (normal-$\uparrow$18.7) 35.8± 2.2 (normal-$\uparrow$8.2) 38.3± 2.2 (normal-$\uparrow$10.9)
     Claude-instant 25.3± 2.9 17.8± 2.2 19.9± 2.4 16.3± 1.7 15.9± 1.6 16.0± 1.6
     Claude-instant + SGCD 39.6± 3.3 (normal-$\uparrow$24.3) 29.6± 2.7 (normal-$\uparrow$11.8) 31.9± 2.5 (normal-$\uparrow$12.0) 37.8± 2.6 (normal-$\uparrow$21.5) 27.2± 2.1 (normal-$\uparrow$11.3) 30.0± 2.1 (normal-$\uparrow$14.0)
     PaLM-2-text 19.1± 1.9 23.0± 2.4 19.9± 2.0 28.1± 2.1 26.7± 2.0 27.1± 2.0
     PaLM-2-text + SGCD 43.5± 3.2 (normal-$\uparrow$24.4) 33.1± 2.7 (normal-$\uparrow$10.1) 35.2± 2.7 (normal-$\uparrow$15.3) 42.9± 2.6 (normal-$\uparrow$23.8) 31.9± 2.1 (normal-$\uparrow$8.9) 34.6± 2.1 (normal-$\uparrow$7.5)
     Average improvement normal-$\uparrow$16.4 normal-$\uparrow$7.9 normal-$\uparrow$10.7 normal-$\uparrow$15.4 normal-$\uparrow$5.8 normal-$\uparrow$8.1
Table 1: Results for closed information extraction, in terms of triplet-based precision, recall, and F1-score (micro-averaged, with bootstrapped 95% confidence intervals) on the Wiki-NRE and SynthIE-text datasets. The results compare the effectiveness of SGCD (blue rows) against two baselines: (1) few-shot-prompted unconstrained decoding with powerful blackbox LLMs (“without logit access”, white rows, Eq. 1) and (2) few-shot-prompted constrained decoding (“CD”) with open-source LLMs (“with logit access”, Eq. 3). Four demonstrations are used in few-shot prompting. LLaMA-7B serves as the constrained generator $P_{\text{cg}}$ for SGCD.
[…]
In our evaluation, we distinguish between sequences that are valid (i.e., that satisfy the constraints) and those that are correct (i.e., those that are equal to the intended output for the given input). A valid output is a prerequisite for being correct, but it is not the sole criterion for correctness.
3.1 Closed information extraction
[…]
Constraints. We use the constraints in Appendix D.2, which restrict entities (1.5 million) and relations (857) to the Wikidata KG, and enforce the structural constraint that outputs must be formatted as sequences of entity–relation–entity triplets.
Datasets and evaluation metrics. We use the Wiki-NRE (Trisedya et al., 2019) and SynthIE-text Josifoski et al. (2023) datasets (details in Appendix D.3). Performance is measured using micro precision, recall, and F1-score.
Results. Table 1 illustrates that powerful blackbox LLMs, even without constrained decoding, already exceed the performance of open-source LLMs that employ constrained decoding. SGCD further enhances the performance of blackbox LLMs. Notably, precision shows a significant increase—by an average of 16.4% on Wiki-NRE and 15.4% on SynthIE-text—, demonstrating the efficacy of constrained decoding in producing outputs that are aligned with the KG. The gains in recall are more modest, with improvements of 7.9% and 5.8% on Wiki-NRE and SynthIE-text respectively, implying that constrained decoding primarily ensures alignment with the KG, rather than completeness. A detailed analysis of output grounding is provided in Appendix D.5.
Degeneration. Our method is built on the assumption that the sketcher will generate a high-quality sketch $y^{*}$, which the constrained generator will refine into $z^{*}$, meeting the specified constraints. Ideally, $z^{*}$ should match or surpass the quality of $y^{*}$, but in certain cases we observe a contrary outcome, which we term “degeneration”. This occurs when the constrained generator diverges from following the sketch $y^{*}$, opting to produce an alternative $z^{*}$ that, while meeting the constraints, is of inferior quality. Note (Table 1) that this issue leads to a (slight) decline in overall performance (in terms of recall) only with GPT-4 as the blackbox LLM, and only on the SynthIE-text dataset.
3.2 Constituency parsing
Penn Treebank
     Method B-precv B-recallv B-F1v Tag accuracyv Validity
With logit access
     LLaMA-2-7B + CD 8.2± 0.1 12.5± 0.2 9.9± 0.1 14.5± 0.1 92.0
     LLaMA-2-13B + CD 12.9± 0.1 20.2± 0.2 15.7± 0.1 16.6± 0.1 95.0
Without logit access
     GPT4 35.0± 0.1 30.6± 0.1 32.7± 0.1 47.0± 0.1 50.3
     GPT4 + SGCD 63.0± 0.1 (normal-$\uparrow$28.0) 57.4± 0.1 (normal-$\uparrow$26.8) 60.1± 0.1 (normal-$\uparrow$27.4) 89.5± 0.1 (normal-$\uparrow$42.5) 100 (normal-$\uparrow$49.7)
     GPT3.5 31.0± 0.1 24.6± 0.1 27.4± 0.1 47.1± 0.1 51.1
     GPT3.5 + SGCD 55.6± 0.1 (normal-$\uparrow$24.6) 44.9± 0.1 (normal-$\uparrow$20.3) 49.7± 0.1 (normal-$\uparrow$22.3) 87.3± 0.1 (normal-$\uparrow$40.2) 100 (normal-$\uparrow$48.9)
     Claude 22.4± 0.2 19.2± 0.1 20.7± 0.1 31.3± 0.1 33.2
     Claude + SGCD 47.4± 0.1 (normal-$\uparrow$25.0) 41.3± 0.1 (normal-$\uparrow$22.1) 44.1± 0.1 (normal-$\uparrow$23.4) 69.0± 0.2 (normal-$\uparrow$37.7) 99 (normal-$\uparrow$65.8)
     Claude-instant 23.6± 0.1 19.1± 0.1 21.1± 0.1 31.1± 0.1 34.6
     Claude-instant + SGCD 54.2± 0.1 (normal-$\uparrow$30.6) 44.7± 0.1 (normal-$\uparrow$25.6) 49.0± 0.1 (normal-$\uparrow$27.9) 77.2± 0.1 (normal-$\uparrow$46.1) 99.5 (normal-$\uparrow$64.9)
     PaLM-2-text 08.8± 0.5 06.7± 0.3 07.6± 0.3 13.2± 0.1 14.0
     PaLM-2-text + SGCD 60.9± 0.1 (normal-$\uparrow$52.1) 48.7± 0.1 (normal-$\uparrow$42.0) 54.1± 0.1 (normal-$\uparrow$46.5) 86.7± 0.1 (normal-$\uparrow$73.5) 100 (normal-$\uparrow$86)
     Average improvement normal-$\uparrow$32.1 normal-$\uparrow$27.4 normal-$\uparrow$29.4 normal-$\uparrow$48.0 normal-$\uparrow$63.1
Table 2: Results for constituency parsing, in terms of bracketing recall, precision, F1-score, and tag accuracy (with bootstrapped 95% confidence intervals), adjusted for validity (B-precv etc.), where each metric is scaled by the percentage of valid parse trees as per the EVALB tool. We compare the same methods as in constituency parsing (Table 1). Few-shot prompting uses three demonstrations.
Task description. Constituency parsing involves breaking down a sentence into its syntactic components to form a parse tree that represents the sentence’s structure. For instance, the sentence “I saw a fox” corresponds to the parse tree [S [NP [PRP I]] [VP [VBD saw] [NP [DT a] [NN fox]]]]. For a visual representation of this tree, see Appendix E, Fig. 5. The instructions $I$ and $I_{\text{cg}}$ are listed in Appendix E.1.
Constraints. We apply the constraints in Appendix E.2 to ensure that parse trees are complete, brackets are balanced, and labels are consistent.
[…]
Results. As indicated in Table 2, even advanced LLMs like GPT-4 face challenges in generating valid parse trees, particularly for longer sentences. LLaMA-2 models (7B, 13B), even using constrained decoding, exhibit much low performance, struggling predominantly with longer inputs Geng et al. (2023). However, with the aid of SGCD, the validity of output parse trees approaches 100% by construction (it may still be slightly less than 100% because the maximum output length is sometimes reached before the end of generation).
The validity adjustment also significantly enhances other metrics, reducing the penalty for invalid outputs. The most common errors in the unconstrained setting include unbalanced brackets and omissions. Appendix E, Fig. 7, provides a detailed example showing how SGCD effectively corrects these errors in draft parse trees.
4 Related work
[…]
Collaborative generation. In the realm of collaborative generation, Vernikos et al. (2023) and Welleck et al. (2023) explored training smaller language models to refine the outputs from larger models for enhanced quality. The skeleton-of-thought approach, proposed by Ning et al. (2023), generates an initial output skeleton and then concurrently develops each segment. Meanwhile, grammar prompting, as suggested by Wang et al. (2023), involves creating a meta-grammar to guide the output of LLMs in producing valid results.
In contrast to these methods, SGCD offers a versatile approach that is not confined to any specific API design. It is particularly aimed at enabling constraint application to the outputs of blackbox LLMs, addressing a distinct challenge in the field.
5 Conclusion
We propose sketch-guided constrained decoding (SGCD), a simple method for constrained decoding with blackbox LLMs that does not require access to next-token logits during generation. Our work is complementary to existing methods for constrained decoding and can be used in conjunction with them. Despite its simplicity, SGCD achieves strong performance on constrained generation tasks, outperforming vanilla LLMs (both blackbox and open-source) by a large margin.
6 Limitations
  • Limitation 1: The degeneration problem sometimes hampers SGCD. More investigation is needed to understand its causes and how to mitigate it.
  • Limitation 2: As LLMs keep getting better, the benefits of SGCD might diminish on some tasks. This is, however, not a limitation of SGCD per se, but rather of constrained decoding in general.
Acknowledgments
We thank Grant Slatton for insightful discussions during the ideation phase. West’s lab is partly supported by grants from Swiss National Science Foundation (200021_185043, TMSGI2_211379), Swiss Data Science Center (P22_08), H2020 (952215), Microsoft Swiss Joint Research Center, and Google, and by generous gifts from Facebook, Google, and Microsoft.
Appendix A Blackbox LLM logit access
Model Logit bias Token probs MMLU
GPT4-0614 Yes Top 5 86.4
GPT3.5-0614 Yes Top 5 70.0
Claude-2.1 No No 78.5
Claude-instant No No 73.4
PaLM-2-text-bison Yes Top 5 78.3
Table 3: The blackbox LLMs we use in our experiments and the access they provide to the logit distribution. MMLU is the mainstream metric for LLM benchmarking.
Logit bias indicates whether the model’s API allows user to pass in a logit bias vector to steer the decoding process, i.e., write access to the logit distribution. Token probs indicates whether the model’s API allows user to access the model’s next token probability distribution, i.e., read access to the logit distribution. MMLU (Hendrycks et al., 2021) is the mainstream metric for LLM benchmarking.
Appendix B Grammar constrained decoding
Grammar-constrained decoding takes a formal grammar $G$ as input and ensures that the output string $w$ is a valid sentence in the formal language $L(G)$ defined by the grammar $G$. This process is achieved through the integration of two key components: a grammar completion engine (Poesia et al., 2022) and a sampling method, e.g. greedy search, nucleus sampling, etc. The grammar completion engine is used to ensure the grammaticality of the output string, while the LLM is used to ensure the plausibility of the output string.
We use Grammatical Framework’s runtime powered completion engine Ranta (2019) with constrained beam search as the sampling method.
Appendix C Logit bias-based iterative decoding
Most blackbox LLM APIs do not provide complete access to the model’s next token probability distribution at each decoding step. Nonetheless, many allow users to input a logit bias parameter to influence the decoding process, i.e., granting users write access but not read access to the model’s logits at each decoding step. This parameter accepts a vector of logits that is added to the logits of the next token probability distribution at each decoding step. By using the logit bias parameter, users can direct the decoding process, effectively masking the logits of invalid tokens. This approach is particularly effective for static constraints, such as lexical constraints (Hokamp and Liu, 2017), where the constraints remain constant throughout the decoding.
However, the logit bias parameter is a static array and does not change during the decoding process. This makes it challenging to apply dynamic constraints, which change as decoding progresses, such as constraints involving membership in formal languages (Deutsch et al., 2019; Poesia et al., 2022; Geng et al., 2023).
A straightforward but costly solution for dynamic constraints is to iteratively invoke the blackbox LLMs API with updated logit bias vectors at each decoding step (Beurer-Kellner et al., 2023; Poesia et al., 2023; Agrawal et al., 2023; Choi et al., 2023). However, this approach is prohibitively expensive. Each API call generates only a single token, and the cost is calculated based on both the input and output tokens [^2] See https://openai.com/pricing for details.. The expense of iteratively calling the blackbox LLMs API with new context and prefix at each step scales quadratically, being $O(n^{2})$ where $n$ is the length of the output sequence. Although methods like those proposed by Beurer-Kellner et al. (2023) and Poesia et al. (2023) use speculation to reduce the number of API calls, the costs can remain high, especially when the constraints are complex.
Appendix D Task 1. closed information extraction
In this section, we provide more details about the closed information extraction task.
D.1 Task instruction
We provide the instruction for the IE task in Figure 2. The few-shot demonstrations are rather long and thus we do not include them here. The full prompt is available in our code repository.
Extract the subject-relation-object triples in fully-expanded format from texts below. The subjects and objects are entities in Wikidata, and the relations are Wikidata properties. Here are a few examples.
(a) Instruction for sketcher
In this task, you will be provided with texts along with draft annotations that represent extracted information triples in the form of subject-relation-object. Your role is to refine these triples to ensure completeness and accuracy. Here are a few examples.
(b) Instruction for constrained generator
Figure 2: Instructions for parsing tasks.
D.2 Grammar
The grammar is defined as follows, where $V$ represents the set of variables, $\Sigma$ the set of terminal symbols, and $P$ the set of production rules:
$V=\{S,T,A,B,C,E,R\},\Sigma=\{\texttt{tokens}\}$
[…]
The outputs are structured as a sequence of triplets, where each triplet is separated by a special marker [e]. Every triplet consists of a subject, a relation, and an object. These elements are each preceded by a special marker: [s] for the subject, [r] for the relation, and [o] for the object, respectively. The subject and object are pre-defined Wikidata entities, and the relation is a pre-defined Wikidata property. This grammar is classified as context-free, more specifically, as a regular grammar.
D.3 IE datasets
The original SynthIE-text and Wiki-NRE datasets comprise 50,000 and 30,000 samples, respectively. To minimize the evaluation cost on Large Language Models (LLMs), we use a smaller subset consisting of 1,000 samples from each dataset.
As noted by Josifoski et al. (2023), the Wiki-NRE dataset displays a significant skew in its relations distribution: the top 10 relations constitute 92% of the triplets, with the top 3 alone accounting for 69%. To ensure our test set accurately reflects the overall dataset, we have downscaled it to 1,000 samples to balance the distribution of relations, as shown in Fig. 3
The SynthIE-text dataset, synthesized by reverse prompting Text-Davinci-003 with triplets from Wikidata, stands out due to its substantial size, diverse content, and high-quality human ratings, as highlighted in Josifoski et al. (2023). This contrasts with prior datasets such as REBEL Huguet Cabot and Navigli (2021), whose annotation quality is low (Josifoski et al., 2022). However, a potential minor bias may exist towards GPT-4 and GPT-3.5, as SynthIE-text was generated from a model in their family, Text-Davinci-003. Despite this, we maintain that this does not compromise the validity of our method, given that our primary focus is on the comparative performance with and without the application of SGCD.
(a) Original relation distribution in WikiNRE test set
(b) Stratified relation distribution in WikiNRE test set
Figure 3: Relation distribution in WikiNRE before and after stratification.
D.4 Discussion of GPT-4 on cIE
An intriguing finding is that SGCD’s performance on the SynthIE-text dataset using GPT-4 (F1=45.6) is marginally lower than that achieved with few-shot prompting alone, without constrained decoding (F1=45.8). Our analysis suggests that the constrained generator occasionally struggles to adhere to the sketcher’s outline, resulting in fewer triplets than expected in the output. This observation is consistent with Wang et al. (2023)’s findings, where constrained decoding was noted to reduce the diversity of the generated samples.
More critically, since SynthIE-text is synthetically generated by reverse prompting Text-Davinci-003 with triplets from Wikidata, its text doesn’t exhibit the naturalness characteristic of the Wiki-NRE dataset. For instance, sentences in SynthIE-text often resemble direct copies with slight alterations from the original entity and relation names. This tendency facilitates the LLMs’ task of grounding entities and relations in the Knowledge Graph (KG), thereby diminishing the necessity for constrained decoding.
However, in real-world scenarios, text is typically more intricate, and grounding entities and relations in the KG is not as straightforward. Despite this, the overall performance enhancement provided by SGCD across various models remains noteworthy, averaging gains of up to 10.7% and 8.1% on Wiki-NRE and SynthIE-text, respectively.
D.5 Grounding analysis
In this study, we delve into the grounding efficacy of GPT-4’s output. A triplet is deemed grounded when both its subject and object entities, as well as the relation, are present in the KG. Furthermore, for a grounded triplet to be considered correct, it must also be part of the target triplet set.
Given that being grounded is essential but not solely adequate for being correct, it is crucial to assess how well GPT-4’s output aligns with the KG. According to the data presented in Table 4, we observe that a significant portion of the output triplets from GPT-4 are not grounded in the KG, amounting to 45% and 37% on the Wiki-NRE and SynthIE-text datasets, respectively. This finding sheds light on the importance of constrained decoding, as it ensures that the output is grounded in the KG, thereby increasing the likelihood of validity.
Ratio of invalid triplets
Wiki-NRE SynthIE-text
Entity Rel Triplet Entity Rel Triplet
     GPT4 19.4 21.9 45.2 7.6 28.1 37.3
     GPT3.5 23.4 50.2 65.8 13.8 52.5 63.3
     Claude 17.0 41.1 55.5 17.4 52.8 64.6
     Claude-ins 19.6 48.4 62.6 13.8 43.3 52.7
     PaLM-2 23.2 55.8 69.1 9.1 56.3 62.4
     Average 20.5 43.5 59.6 12.3 46.6 56.1
     SGCD 0 0 0 0 0 0
Table 4: Triplets grounding analysis. We report the percentage of generated entities, relations, and triplets that are not present in the knowledge catalogue in few-shot unconstrained setting. The grounding precision for constrained methods is 100% by construction, and thus 0% invalid triplets.
Appendix E Task 2. constituency parsing
In this section, we provide more details about the constituency parsing task.
E.1 Task instruction
We provide the instruction for the CP task in Figure 4. The few-shot demonstrations are rather long and thus we do not include them here. The full prompt is available in our code repository.
Perform constituency parsing on the provided sentences in accordance with the Penn TreeBank annotation guidelines. Here are a few examples.
(a) Instruction for sketcher
In this task, you will be provided with a draft annotations that represent the parse tree of a sentence in Penn TreeBank format. Your task is to rewrite the parse tree and fix error if any. Here are a few examples.
(b) Instruction for constrained generator
Figure 4: Instructions for parsing tasks.
E.2 Constraints and grammar
[…]
[.S [.NP [.PRP I ] ] [.VP [.VBD saw ] [.NP [.DT a ] [.NN fox ] ] ] ]
(a) The correct constituency parse tree
[…]
[.S [.VP [.VBD I ] [.NP [.DT saw ] [.NN a ] [.NN fox ] ] ] ]
(b) A grammatical but incorrect parse tree
Figure 5: Parse trees for the sentence “I saw a fox”.
Here we describe the grammar used to constrain the generative constituency parsing task.
Linearization. A constituency parse tree is inherently a recursive structure. To effectively represent this tree as a sequence of tokens for generation by a Large Language Model (LLM), a linearization is required. Two common strategies for this linearization are pre-order traversal and post-order traversal.
We have chosen to adopt the pre-order traversal strategy. This approach is also the default method used in the PYEVALB tool Sekine and Collins (2008) and in the construction of the Penn Treebank Marcus et al. (1993). As an illustration, the parse tree in Fig. 4(a) is linearized in the following format: [S [NP [PRP I]] [VP [VBD saw] [NP [DT a] [NN fox]]]].
The linearised parse tree needs to satisfy the following structural constraints:
  • Completeness: Every word in the sentence needs to be included in the parse tree.
  • Balanced brackets: At any point in the linearized parse tree, the right bracket ] should close a previously unclosed left bracket [ and every left bracket [ should be eventually closed by a right bracket ].
  • Label consistency: The label of terminal and non-terminal nodes needs to be consistent with the Penn Treebank format.
Grammar. The tree structure of the parse tree is usually captured by a context-free grammar. However, the context-free grammar is not sufficient to capture the completeness constraint, motivating the use of a more restrictive grammar. We use the same grammar as in Geng et al. (2023) as shown in Fig. 6.
$S\to B_{0,0}$
$B_{i,j}\to[\alpha\,(B_{i,j+1}\mid C_{i,j+1})];$
$C_{i,j}\to x_{i}\,(C_{i+1,j}\mid E_{i+1,j});$
$C_{n,j}\to E_{n,j};$
$E_{i,j+1}\to]\,(E_{i,j}\mid B_{i,j});$
$E_{n,j+1}\to]\,E_{n,j};$
$E_{n,0}\to\varepsilon;$
$\text{where}\,\alpha=(S\mid NP\mid VP\mid\ldots)$ $\text{and}\,x_{i}\in\text{tokens}$
Figure 6: Grammar for constituency parsing used in constrained decoding.
The grammar falls into the category of regular grammar and is input-dependent. it reproduces the input sentence, represented as a sequence $x=\langle x_{0},\dots,x_{n-1}\rangle$ of words, in left-to-right order, interspersing it with node labels and balanced brackets. In order to guarantee balanced brackets, the non-terminals $B_{i,j}$ count the number of opened left brackets [ using the second subscript index $j$, and the rules ensure that the number of closed brackets can never exceed the number of previously opened brackets.
E.3 Example of constituency parsing correction
We provide an example to show the correction from our method in Figure 7. As shown in the figure, the parse tree of the input sequence is not valid because the number of closing brackets is 2 times less than the number of opening brackets and some tags are not consistent with the Penn Treebank format. The constrained generator generates a sequence of tokens that is consistent with the grammar and fixes the parse tree to be valid.
However, the fixed parse tree is not yet identical to the target parse tree, cf. blue nodes in Figure 7. This is expected because the constrained generator can only fix the parse tree to be consistent with the grammar but cannot guarantee that the parse tree is identical to the target parse tree.
Figure 7: Example to show the correction process of our method. The red nodes are the nodes that are corrected by constrained generator. The blue nodes are the nodes that different from the target parse tree( some of them are not corrected by constrained generator because they are not violating the grammar).
E.4 Validity adjusted metric
We use the validity adjusted metric as done in Hemmer et al. (2023) to measure the effectiveness of our method. It is defined as follows:
$\text{Metric}^{\text{v}}=\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}(\text{valid}(y_{i}))\cdot\text{Metric}(y_{i})$
$\text{valid}(y_{i})=\begin{cases}1&\text{if the parse tree is valid}\\ 0&\text{otherwise}\end{cases}$
Appendix F Data contamination risk
There is a rising concern over the data contamination risk of evaluating LLMs on downstream tasks. The datasets of our experiments are publicly available on internet so there is a risk that the models may have seen the data, such as the ground true parse tree of Penn Treebank during pretraining. However, the risk of data contamination is independent of our method and doesn’t affect the validity of our conclusions.